An ocean fishing boat is drifting just above a school of tuna on afoggy day. Without warning, an engine backfire occurs on anotherboat 1.0 km away. How much time elapses before the backfire isheard (a) by the fish, and (b) by the fishermen?
Answer #1
Distance between fire & boat d = 1 km =1000m
velocity of sound in air v = 343 m/s
velocity of sound in water v w = 1560m/s
a) time taken to rach the sound to fish   t =d / v
                                                           t = 1000 / 1560
                                                            = 0.64 s
b) time taken to rach the sound to fisherman  t = d / v
                                                                         =1000/ 343
                                                                        = 2.9s